QUESTION 11 A Find the singular values of the matrix. - 4 0 02 O A 16,-2. OB 16,4 Oc 4,2 OD.-16,4 OE-4,2
Answer #1

Singular\ value\ of\ a\ matrix\ A\ are\ square\ root\ of\ eigen\ values\ of\ A^TA.\\ here,\ A=\begin{bmatrix} -4& 0\\ 0& 2 \end{bmatrix},\\ A^TA=\begin{bmatrix} 16&0 \\ 0 & 4 \end{bmatrix},\\ Eigen\ values\ of\ A^TA\ are\ 16,4.\\ square\ root\ of\ eigen\ values\ are\ 4,2.\\ hence,\ singular\ values\ are\ 4,2.